Let ƒ(x) = 3x2 – 2x and g(x) = x – 6. Find each value.
THANKS FOR YOUR HELP.. I APPRICATE IT.:)

1. (f-g)(2)

2. (f/g)

3. (gof)(3)

4. (fog)(0)

Answer by stanbon(48558)

Let ƒ(x) = 3x2 – 2x and g(x) = x – 6. Find each value.
THANKS FOR YOUR HELP.. I APPRICATE IT.:)
1. (f-g)(2)
(f-g) = 3x^2-2x -(x-6) = 3x^2-3x+6
(f-g)(2) = 3(2^2)-3*2+6 = 3*4 = 12
---------------------------------------
2. (f/g)
(f/g) = (3x^2-2x)/(x-6)
---------------------------------------

3. (gof)(3) = g[f(2)] = g[3*2^2-2*2] = g[8] = 8-6 = 2
-----------------------------------------------------------
4. (fog)(0) = f[g(0)] = f[-6] = 3(-6)^2-2(-6) = 108 +12= 120
===================================================================
Cheers,
Stan H.